This afternoon I tuned the handoff script for the SRC, after that Rob eralier during the day had already adjusted that for PRC. To do that, I followed the procedure in the Wiki.
After that the SRC could get locked with the double demod signals. the open loop transfer function emasurement on the PRC loop showed that it was nearly unstable. Rob reduced a little its gain to improve the stability.
The DD handoff is now working and we can get back to locking the interferometer.